Can someone help me with these problems?
Write an equation of the line containing the given point and perpendicular to the given line. Express your answer in the for form of y=mx +b
(6,7); 3x + y=8
_______________________________________________________
Find the slope, if it exists, of the line containing the pares of points.
(-17.7,13.5)and (-17.5, 8.7)

Answer by stanbon(75887)

Write an equation of the line containing the given point and perpendicular to the given line. Express your answer in the for form of y=mx +b
(6,7); 3x + y=8
---
The given line has slope-intercept form y = -3x+8
So it's slope is -3.
The equation you want must have slope 1/3 if it is going to a perpendicular line.
---
So find the equation of a line with slope 1/3 passing thru (6,7)
y = mx + b
7 = (1/3)6 + b
21/3 = 6/3 + b
b = 15/3 = 5
---
Equation:
y = (1/3)x+5
===================
_______________________________________________________
Find the slope, if it exists, of the line containing the pairs of points.
(-17.7,13.5)and (-17.5, 8.7)
----
Slope = (8.7-13.5)/(-17.5--17.5)
Since the denominator is zero the slope is "undefined".
===================
If you plot the two points and draw a line thru them
you will see the line is vertical and passes thru
all points where x = -17.5.
===================
Cheers,
Stan H.
===========